OpenStudy (anonymous):

Slope Intercept form 5x-9y=-7

OpenStudy (apoorvk):

slope intercept form is the form y=mx+c so re-configure your equation to that form. first divide by 5, then rearrange.

OpenStudy (anonymous):

slope is -a\b u should first re-arange the equation,so it will be 5x-9y+7=0 the accoring to the formula it gonna be ax+by-c=0

OpenStudy (anonymous):

in your equation : 5x-9y+7.. a=5,b=-9 ,so -a\b=-5\-9=5\9 ,hope u understand it :)